OpenStudy (anonymous):

Which unit rate is equivalent to 19 miles per gallon? three gallons over fifty seven miles fifty seven miles over three gallons thirty seven miles over four gallons four gallons over thirty seven miles

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Michele_Laino r u there?

OpenStudy (michele_laino):

As I wrote before, I think the second option, because: \[\frac{ 57 }{ 3 }=\frac{ 19 }{ 1 }\]

OpenStudy (michele_laino):

@77777jeannie77777

OpenStudy (anonymous):

yes.. oh now i understand

OpenStudy (anonymous):

An oil company fills 1 over 12 of a tank in 1 over 4 hour with gasoline. At this rate, which expression can be used to determine how long will it take for the tank to fill completely? 1 over 4 ⋅ 1 over 12 1 over 12 ⋅ 4 1 over 4 ⋅ 12 12 ⋅ 4

OpenStudy (anonymous):

An oil company fills 1 over 12 of a tank in 1 over 4 hour with gasoline. At this rate, which expression can be used to determine how long will it take for the tank to fill completely? 1 over 4 * 1 over 12 1 over 12 * 4 1 over 4 * 12 12 * 4

OpenStudy (anonymous):

what do you think @Michele_Laino?

OpenStudy (michele_laino):

I think the second option, because, you have to calculate: \[\frac{ \frac{ 1 }{ 12 } }{ \frac{ 1 }{ 4 } }=...\] please continue

OpenStudy (michele_laino):

please the first fraction multiply the inverse of the second fraction

OpenStudy (anonymous):

so then multiply 1/12 over 1/4 * 1/4 over 1/12?

OpenStudy (michele_laino):

oops sorry ...multiplies...

OpenStudy (anonymous):

so then i was right?

OpenStudy (michele_laino):

\[\frac{ 1 }{ 12 }*\frac{ 4 }{ 1 }\]

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

its 4/12

OpenStudy (michele_laino):

better is \[\frac{ 1 }{ 12 }*4\] so the second option

OpenStudy (anonymous):

ok, thx :D

OpenStudy (michele_laino):

thanks!

OpenStudy (michele_laino):

please, I have to go to have dinner! wait a moment please!